A space probe in the shape of the sphere x ^ 2 + y ^ 2 + z ^ 2 = 30 enters Earth's atmosphere and its surface begins to heat. After 1 hour, the temperature at the point (z. y. :) on the probe's surface is T(x, y, z) = x - 2y + 5z Find the hottest point on the probe's surface.


1
Expert's answer
2021-11-11T18:43:01-0500

Maximize T(x,y,z)=x-2y+5z

Subject to g(x,y,z)=x2+y2+z2-30=0

"L=T-\lambda g"

"=x-2y+5z-\lambda(x^2+y^2+z^2-30)"

"L_x=1-2\lambda x=0 \implies2\lambda x=1\implies x=\frac{1}{2\lambda}"

"L_y=-2-2\lambda y=0\implies 2\lambda y=-2\implies y=\frac{-1}{\lambda}"

"L_z=5-2z\lambda=0\implies 2\lambda z=5\implies z=\frac{5}{2\lambda}"

"g(x,y,z)=(\frac{1}{2\lambda})^2+(\frac{-1}{2\lambda})^2+(\frac{5}{2\lambda})^2-30=0"

"\frac{1}{4\lambda^2}+\frac{1}{\lambda^2}+\frac{25}{4\lambda^2}=30"

"\frac{1+4+25}{4\lambda^2}=30"

"\frac{29}{4\lambda^2}=30"

"4\lambda^2*30=29"

"\lambda^2=\frac{29}{4*30}"Β ="\frac{29}{120}"

"\lambda=\sqrt{\frac{29}{120}}="Β 0.4916

"\lambda=0.4916, x=\frac{1}{2\lambda}=1.0171,y=\frac{-1}{\lambda}=-2.0342,z=\frac{5}{2\lambda}=5.0854"

"T(1.0171,-2.0342,5.0854)=(1.0171-2(-2.0342)+5(5.0854)"

=30.5125

Hottest point on the probe surface is (1.0171,-2.0342,5.0854)